Peter invests in a public project which needs a $500,000 payment now and a second $900,000 payment 10 years from now. This project is considered with a perpetual life. The effective annual interest rate is 12%. What is the approximate capitalized cost?

A.

$289,800

B.

$786,740

C.

$789,800

D.

$5,585,180
Answer #1

Present value of cash outflow at year 0 =500,000

Present value of cash flow at year 10

=PV(rate,nper,pmt,fv)

=PV(12%,10,,900000) =289776

Total of capitalized cost =500,000+289,776 =789,776

The answer is C)

$789,800